OpenStudy (anonymous):

What is the equation, in point-slope form, of the line that has a slope of 6 and passes through the point (–1, –8)?

OpenStudy (gabylovesyou):

Point slope form is y - y1 = m(x - x1) All you have to do is substitute the letters for the numbers they gave you.... we dont know what y is so you leave it as y... y1 is -8 so you already have y - (-8). m is the slope... so 6 .. we have y - (-8) = 6 we dont have x.... so y - (-8) = 6(x - and x1 is -1 so the equation is y - (-8) = 6(x - (-1)) :)